# TDS-CLIP: Temporal Difference Side Network for Image-to-Video Transfer Learning

### v1

@@ Conclusion
In this paper, we introduce a novel approach to VAR task that balances memory efficiency with effective temporal modeling. Our method addresses the challenging task of transferring large, task-specific vision-language models to the domain of video action recognition, which can be facilitated for researchers with limited hardware. The core innovation of our work is the development of a temporal modeling adapter tailored for side networks and the integration of spatio-temporal features, all while circumventing the need for backpropagation through the CLIP pre-trained model. Comprehensive experiments conducted on various benchmark datasets demonstrate that our method significantly enhances temporal modeling capabilities. In the era of large language models dominated by Transformers structures, we hope our work will inspire further exploration of parameter-efficient fine-tuning methods for video temporal modeling.

### v2

@@ V Conclusion
In this paper, we introduce a novel approach to video action recognition task that balances memory efficiency with effective temporal modeling. Our method addresses the challenging task of transferring large, task-specific vision-language models to the domain of video action recognition, which can be facilitated for researchers with limited hardware. The core innovation of our work is the development of a temporal modeling adapter tailored for side networks and the integration of spatio-temporal features, all while circumventing the need for backpropagation through the CLIP pre-trained model. Comprehensive experiments carried out on various benchmark datasets demonstrate that our method significantly enhances the temporal modeling capabilities. The disadvantage of this approach is that it sacrifices zero-shot generalization capability. In the era of large language models dominated by Transformers structures, we hope our work will inspire further exploration of parameter-efficient fine-tuning methods for video temporal analysis.